OpenStudy (anonymous):

Solve the system: x^2+y^2=16 x+y=0

OpenStudy (agreene):

\[(-2\sqrt{2},2\sqrt{2});(2\sqrt{2},-2\sqrt{2})\]

OpenStudy (agreene):

those pair of equations create the null sign... I approve.

OpenStudy (anonymous):

thank you so much